# Scaffold-Aware Generative Augmentation and Reranking for Enhanced Virtual Screening

### v1

@@ 5 Limitations and Future Work
Despite ScaffAug’s promising performance, some limitations warrant discussion. First, while our generative approach produces chemically valid molecules, it does not guarantee synthetic accessibility or practical feasibility. This gap between computational generation and practical synthesis represents a significant challenge when applying ScaffAug for actual drug discovery campaigns. Second, as a ligand-based virtual screening approach, ScaffAug neglects protein target structures, potentially reducing effectiveness compared to structure-based methods. Future work could address this limitation by integrating protein structure into the generative process, leveraging protein-aware diffusion models(Nobrega et al., 2022; Schneuing et al., 2022) to generate synthetic compounds with a greater likelihood of target activity. Such extensions would enable the framework to benefit from both ligand-based efficiency and structure-based accuracy, potentially yielding even more powerful virtual screening capabilities. Additionally, while our current approach utilizes pseudo-labeling for safely integrating synthetic data, alternative semi-supervised learning methods could be explored to combine unlabeled synthetic data with labeled examples more effectively.
@@ 6 Conclusion
In this work, we introduced ScaffAug, a novel scaffold-aware framework that addresses three major challenges in virtual screening: class imbalance, structural imbalance among active molecules, and limited scaffold diversity in top predictions. Our augmentation module uses scaffold-aware sampling and graph diffusion models to generate synthetic molecules that preserve critical scaffolds, the self-training module safely incorporates these molecules through confidence-based pseudo-labeling, and the reranking module increases scaffold diversity while maintaining hit identification performance. Extensive experiments across five therapeutic target classes demonstrate that ScaffAug consistently outperforms existing baselines on multiple evaluation metrics for both random and scaffold splits. ScaffAug represents a significant advancement in computational drug discovery by leveraging generative AI to enhance virtual screening, particularly in early-stage campaigns where identifying structurally diverse active compounds is critical.
@@ A.2 Baselines
DCT (Data-Centric Transfer) uses a diffusion model pretrained on large unlabeled graph collections to learn the graph distribution, then performs task-conditioned augmentation: for each labeled example, it perturbs the graph via forward diffusion and then reverses it under dual objectives of preserving the original label and promoting structural diversity (Liu et al., 2023a). Augmented graphs thus carry minimal sufficient knowledge from the unlabeled domain and enrich the training set.
@@ A.3 Data Splits
For scaffold-based splits, we propose a benchmark standard that explicitly supports scaffold hopping, a central objective in drug discovery for generating structurally novel compounds with improved properties and patentability Hu et al. (2017). We use Bemis–Murcko (BM) scaffolds Bemis & Murcko (1996) with a 3:1:1 training:validation:test ratio, assigning any scaffold bin exceeding 10% of the dataset to the training set. This ensures scaffold diversity is preserved across all splits.
@@ A.4 Evaluation Metrics Details
The ranged logAUC metric (Mysinger & Shoichet (2010)) captures performance in the low false-positive regime, emphasizing the left-hand portion of the ROC curve. Since the optimal threshold is unknown, the metric integrates the area under the ROC curve across a restricted FPR interval. Taking the logarithm accentuates contributions from extremely small FPR values. Following prior studies (Liu et al. (2023b); Golkov et al. (2020)), we adopt the interval [0.001, 0.1]. A perfect predictor achieves logAUC, while random guessing gives a value of approximately 0.0215, computed as:
@@
BEDROC (Pearlman & Charifson (2001)) ranges between 0 and 1, quantifying a model’s ability to prioritize active molecules early in a ranked list. It is derived from the Robust Initial Enhancement (), normalized using its theoretical minimum (all actives at the end of the list) and maximum (all actives at the top). The definitions are:
@@
Discounted Cumulative Gain (Järvelin & Kekäläinen (2017)) measures ranking quality by penalizing relevant items that appear deeper in the list. First, the simpler Cumulative Gain (CG) counts the actives in the top 100:

### v2

@@ 5 Limitations and Future Work
Despite ScaffAug’s promising performance, some limitations warrant discussion. First, while our generative approach produces chemically valid molecules, it does not guarantee synthetic accessibility or practical feasibility. This gap between computational generation and practical synthesis represents a significant challenge when applying ScaffAug for actual drug discovery campaigns. Second, as a ligand-based virtual screening approach, ScaffAug neglects protein target structures, potentially reducing effectiveness compared to structure-based methods. Future work could address this limitation by integrating protein structure into the generative process, leveraging protein-aware diffusion models(Schneuing et al., 2024) to generate synthetic compounds with a greater likelihood of target activity. Such extensions would enable the framework to benefit from both ligand-based efficiency and structure-based accuracy, potentially yielding even more powerful virtual screening capabilities. Additionally, while our current approach utilizes pseudo-labeling for safely integrating synthetic data, alternative semi-supervised learning methods could be explored to combine unlabeled synthetic data with labeled examples more effectively.
@@ 6 Conclusion
In this work, we introduced ScaffAug, a novel scaffold-aware framework that addresses three major challenges in virtual screening: class imbalance, structural imbalance among active molecules, and limited scaffold diversity in top predictions. Our augmentation module uses scaffold-aware sampling and graph diffusion models to generate synthetic molecules that preserve critical scaffolds, the self-training module safely incorporates these molecules through confidence-based pseudo-labeling, and the reranking module increases scaffold diversity while maintaining hit identification performance. Extensive experiments across five therapeutic target classes demonstrate that ScaffAug consistently outperforms existing baselines on multiple evaluation metrics for both random and scaffold splits. ScaffAug represents a significant advancement in computational drug discovery by leveraging generative models to enhance virtual screening, particularly in early-stage campaigns where identifying structurally diverse active compounds is critical.
@@ A.1 Dataset Details
The WelQrateLiu et al. (2024b) collection establishes a high-quality benchmark through the rigorous curation of experimental data sourced from PubChem. The suite comprises nine distinct datasets that encompass a variety of critical drug targets, all characterized by the challenging, low active-to-inactive ratios typical of realistic drug discovery scenarios.
@@ A.2 Baselines
DCT Liu et al. (2023a) uses a diffusion model pretrained on large unlabeled graph collections to learn the graph distribution, then performs task-conditioned augmentation: for each labeled example, it perturbs the graph via forward diffusion and then reverses it under dual objectives of preserving the original label and promoting structural diversity. Augmented graphs thus carry minimal sufficient knowledge from the unlabeled domain and enrich the training set.
InstructMol
InstructMol (Cao et al., 2025) introduces a flexible semi-supervised learning framework for molecular property prediction that splits the workflow into two models: a target model that assigns pseudo-labels to large unlabeled molecular datasets, and an instructor model that estimates the reliability (confidence scores) of those pseudo-labels. The target model is then trained on both labeled and pseudo-labeled data using a loss re-weighting strategy guided by the instructor’s confidence scores, enabling effective utilization of unlabeled data without relying on domain transfer between disjoint pre-training and fine-tuning stages.
@@ A.3 Data Splits
For scaffold-based splits, we propose a benchmark standard that explicitly supports scaffold hopping, a central objective in drug discovery for generating structurally novel compounds with improved properties and patentability Hu et al. (2017). We use Bemis–Murcko (BM) scaffolds Bemis and Murcko (1996) with a 3:1:1 training:validation:test ratio, assigning any scaffold bin exceeding 10% of the dataset to the training set. This ensures scaffold diversity is preserved across all splits.
@@ A.4 Evaluation Metrics Details
The ranged logAUC metric (Mysinger and Shoichet (2010)) captures performance in the low false-positive regime, emphasizing the left-hand portion of the ROC curve. Since the optimal threshold is unknown, the metric integrates the area under the ROC curve across a restricted FPR interval. Taking the logarithm accentuates contributions from extremely small FPR values. Following prior studies (Liu et al. (2023b); Golkov et al. (2020)), we adopt the interval [0.001, 0.1]. A perfect predictor achieves logAUC, while random guessing gives a value of approximately 0.0215, computed as:
@@
BEDROC (Pearlman and Charifson (2001)) ranges between 0 and 1, quantifying a model’s ability to prioritize active molecules early in a ranked list. It is derived from the Robust Initial Enhancement (), normalized using its theoretical minimum (all actives at the end of the list) and maximum (all actives at the top). The definitions are:
@@
Discounted Cumulative Gain (Järvelin and Kekäläinen (2017)) measures ranking quality by penalizing relevant items that appear deeper in the list. First, the simpler Cumulative Gain (CG) counts the actives in the top 100:
@@ A.6 Self-Training Algorithm
A.7 Ablation Study for Self-Training Module
In this section, we demonstrate the effectiveness of the self-training module. We evaluate a variant, , which assigns the active label to all generated molecules and merges them directly with the original dataset. Using a GCN backbone, we observe that this direct augmentation strategy performs worse than naive GCN training. Conversely, the proposed self-training module consistently outperforms the GCN baseline across all five datasets under both random and scaffold splits. Note that we apply oversampling of active molecules during training to all experiment settings because of the inherent low active rate of the WelQrate datasets.
| AID1798 | AID2689 | AID463087 | AID485290 | AID488997 | ||||||
| logAUC | ||||||||||
| GCN | 0.1080.022 | 0.0560.017 | 0.2370.028 | 0.1600.024 | 0.3030.013 | 0.2650.019 | 0.1670.006 | 0.1640.010 | 0.1640.031 | 0.1150.035 |
| 0.0950.009 | 0.0600.009 | 0.1380.012 | 0.1330.015 | 0.2850.008 | 0.2640.007 | 0.0980.008 | 0.0940.004 | 0.1510.016 | 0.1410.010 | |
| 0.1090.018 | 0.0430.017 | 0.2720.031 | 0.1690.017 | 0.3560.014 | 0.2930.019 | 0.1740.011 | 0.1660.020 | 0.2370.033 | 0.1080.014 | |
| BEDROC | ||||||||||
| GCN | 0.2020.029 | 0.1180.021 | 0.3700.033 | 0.2700.030 | 0.4740.013 | 0.4340.021 | 0.2740.011 | 0.2700.012 | 0.2470.038 | 0.1990.046 |
| 0.1840.013 | 0.1320.012 | 0.2540.019 | 0.2480.022 | 0.4490.009 | 0.4290.008 | 0.1810.010 | 0.1800.007 | 0.2250.027 | 0.2250.009 | |
| 0.1980.020 | 0.0970.029 | 0.4030.036 | 0.2880.024 | 0.5090.014 | 0.4540.021 | 0.2840.012 | 0.2730.019 | 0.3260.035 | 0.1880.018 | |
| GCN | 7.6723.918 | 3.2292.322 | 38.90516.047 | 15.70821.619 | 30.3173.143 | 24.6514.570 | 23.6825.255 | 24.1316.709 | 34.29721.258 | 13.40413.871 |
| 6.0703.665 | 1.8171.388 | 25.3739.265 | 12.99610.394 | 28.7461.899 | 25.2881.929 | 10.5575.109 | 9.7194.226 | 38.40711.413 | 19.85314.100 | |
| 7.7563.918 | 2.0271.363 | 52.43727.012 | 18.69013.254 | 41.9354.543 | 30.9023.786 | 23.34813.917 | 23.25915.113 | 66.93923.045 | 11.6234.853 | |
| GCN | 0.4630.207 | 0.1720.148 | 0.3040.107 | 0.1300.178 | 5.3320.596 | 3.6660.605 | 1.2980.265 | 1.2020.248 | 0.6710.522 | 0.2110.212 |
| 0.3770.183 | 0.1540.111 | 0.2280.094 | 0.0830.079 | 5.5970.443 | 4.4310.235 | 0.5970.254 | 0.5260.313 | 0.7010.260 | 0.2980.200 | |
| 0.4790.202 | 0.0810.053 | 0.4900.331 | 0.1250.082 | 7.3270.625 | 5.2030.642 | 1.2090.725 | 1.0840.488 | 1.2860.660 | 0.2130.093 | |
A.8 Visualization of Augmentation Module
A.8.1 Generated Scaffolds
In the figure below, we show the sampled active scaffold used for augmentation. The remaining scaffolds are novel and do not exist in the original active space of the training set. They are ranked by their similarity scores relative to the sampled scaffold. Although the scaffold is fixed during scaffold-conditioned generation, we can still increase scaffold diversity across the generated active molecules by editing this core structure.
A.8.2 Scaffold Extension Steps
Below, we visualize the scaffold-conditioned generation process, corresponding to the reverse step described in Alg. 3. We use a maximum of 50 reverse steps. Initially, the unknown region consists of noise sampled from the unlabeled data distribution. As the reverse steps progress, this noise gradually transforms into meaningful substructures around the scaffold, eventually forming valid molecules for augmentation.